# CREA: A Collaborative Multi-Agent Framework for Creative Content Generation with Diffusion Models

### v1

@@ D.1.3 Self-Enhancement with Optional User-Guidance
D.2 Creative Image Editing
@@
Creative image editing in CREA follows the same multi-agent conversational structure as the generation pipeline but is adapted to operate on a user-provided input image. Instead of synthesizing a new image from scratch, the Generative Executor (A3) applies disentangled edits using models like ControlNet, guided by a high-creativity prompt crafted during the pre-generation planning phase. The agents—Creative Director, Prompt Architect, Art Critic, and Refinement Strategist—engage in the same collaborative process of blueprint creation, evaluation, and refinement. However, the Executor now conditions on the input image to preserve key visual elements while transforming its aesthetic, structure, or narrative creatively. Post-generation evaluation and self-enhancement phases remain identical, ensuring that the final edited image not only retains its semantic core but also meets the creativity threshold defined by the agent team.
Appendix E Creative Video Generation
To demonstrate the versatility and extensibility of our agentic framework, we extend CREA to creative video generation. We conduct experiments using the CogVideoX [74] model as the generative backbone. This evaluation serves as a proof of concept for applying our multi-agent creativity principles beyond static imagery.
Instead of initializing the process with a static creative blueprint, our Creative Director agent generates a structured creative video plan containing key fields: Subject, Action, Setting, Style, and optional Additional Details. This plan is passed to the Prompt Architect, who composes a coherent and high-creativity video prompt via contrastive prompt fusion, similar to our image generation pipeline. The Generative Executor then synthesizes the video using CogVideoX.
As shown in Figure 8, our method produces significantly more imaginative and visually engaging results compared to baseline prompts such as “a creative object.”
Appendix F Additional Qualitative Results
We provide various qualitative results for both generation and editing to demonstrate our method’s ability to produce both diverse and highly creative images. Please see Figures 17-40.

### v2

@@ C.1.3 Self-Enhancement with Optional User-Guidance
C.2 Creative Image Editing
@@
Creative image editing in CREA follows the same multi-agent conversational structure as the generation pipeline but is adapted to operate on a user-provided input image.Instead of synthesizing a new image from scratch, the Generative Executor (A3) applies disentangled edits using models like ControlNet, guided by a high-creativity prompt crafted during the pre-generation planning phase.The agents- Creative Director, Prompt Architect, Art Critic, and Refinement Strategist engage in the same collaborative process of blueprint creation, evaluation, and refinement.However, the Executor now conditions on the input image to preserve key visual elements while transforming its aesthetic, structure, or narrative creatively.Post-generation evaluation and self-enhancement phases remain identical, ensuring that the final edited image not only retains its semantic core but also meets the creativity threshold defined by the agent team.
Appendix D Additional Experiments
D.1 Additional Ablations
Please see Fig.16 for an additional ablation on how providing a negative prompt (ie “A normal <obj>") affects generation using SDXL.
D.2 User Study Details
Please see Fig.17 shows a screenshot from our user study.
D.3 Additional Quantitative Results
Creative Editing: In order to isolate the impact of our agentic framework from potential improvements due to the choice of base model (e.g., Flux-1-dev), we conducted a controlled comparison with RF-Inversion [rout2024semantic], a recent Flux-based image editing method.As shown in Table 4, CREA significantly outperforms RF-Inversion across all quantitative metrics, despite both using the same underlying image generator, supporting our claim that the performance gains stem from our structured prompt design and agentic reasoning architecture, rather than from the base model alone.
| Method | CLIP | LPIPS* | Vendi | DINO | FID | KID |
| RF-Inversion | 0.369 ± 0.045 | 0.269 ± 0.084 | 3.35 ± 1.53 | 0.743 ± 0.179 | 309.73 | 21.76 |
| Ours (CREA) | 0.417 ± 0.030 | 0.414 ± 0.157 | 3.70 ± 1.97 | 0.744 ± 0.185 | 294.19 | 14.02 |
Creative Generation: To further differentiate CREA from prior work, we conduct a targeted human evaluation comparing our framework with GenArtist [wang2024genartist], a recent single-agent system for general multimodal editing.While GenArtist was not designed specifically for creative editing, we adapt it by including the word “creative” in the editing prompt (e.g., “a creative bike”) to provide a fair comparison. Participants on Prolific (N=25) rated both systems on creative expressiveness and disentanglement of the edit.As shown Table 5, CREA significantly outperforms GenArtist on both axes, suggesting that our multi-agent, creativity-grounded framework enables more interpretable and compositionally novel results.
| Method | Q1: Creativity | Q2: Disentangled Edit |
| GenArtist | 2.893 ± 1.206 | 2.850 ± 1.193 |
| Ours (CREA) | 4.299 ± 0.903 | 3.726 ± 1.017 |
D.4 Additional Qualitative Results
D.4.1 Editing and Generation Examples
D.4.2 Fine-Grained Edits
As shown in Fig.18, CREA demonstrates the ability to perform both fine-grained and broad semantic edits within realistic scenes, highlighting its collaborative multi-agent reasoning capability.Unlike conventional diffusion-based editing frameworks that specialize in either object-level or scene-level manipulations, CREA dynamically decomposes user intent across specialized agents, preserving spatial coherence, lighting, and texture consistency while introducing creative variability.This allows CREA to seamlessly transition from minor contextual additions (e.g., inserting small objects) to global transformations (e.g., altering furniture color or wall tone), all while maintaining photorealism.
D.4.3 GPT-4o as Both Agent and Image Generator
We further demonstrate CREA’s multi-agent creativity when integrated with GPT-4o as both the reasoning agent and the image generator.In this configuration, GPT-4o collaboratively interprets user prompts, decomposes them into semantic sub-goals, and synthesizes visually coherent outcomes using diffusion-based rendering.The results (See Fig.19) illustrate CREA’s dual capability across creative image editing (left) and creative image generation (right).In editing tasks, CREA enhances existing inputs with imaginative, stylistically rich augmentations while preserving structural fidelity.In generation tasks, it transforms minimal textual cues into surreal, high-fidelity visuals that balance novelty with realism, showcasing the system’s ability to produce conceptually grounded yet artistically diverse outputs.